Subject: APO 3.5m users committee minutes, Dec 2012

From: Suzanne Hawley

Submitted: Thu, 6 Dec 2012 20:24:11 -0800 (PST)

Message number: 1298 (previous: 1297, next: 1299 up: Index)

 		APO 3.5-m Users Committee Phonecon, 12/03/12

Attending:  Suzanne Hawley, Mark Klaene, John Bally, Scott Anderson, Jon 
Holtzman, Rene Walterbos, Michael Strauss, Mary Beth Kaiser, Al Harper, 
Remy Indebetouw, Bill Ketzeback, and Bruce Gillespie.

Absent:  none


User feedback and comments from institutional representatives:

o Washington (Scott Anderson) - Scott mentioned that John Wisniewski has 
had recent experience using the echelle and reports that the flip mirror 
is now working well.  He has not seen any increase in overall throughput 
since the new grating was installed.  Mary Beth added that JHU users have 
seen a drop (down about 25%) in the echelle lamp cals.  Suzanne and Bill 
noted that this is normal after a lamp changeout.

o Chicago (Al Harper) - Al had nothing new to report.

o New Mexico State (Jon Holtzman) - Jon had nothing new to report.

o Princeton (Michael Strauss) - Michael had nothing new to report.

o Colorado (John Bally) - John had nothing new to report.

o Virginia (Remy Indebetouw) - Remy had nothing new to report.

o Johns Hopkins (Mary Beth Kaiser) - Mary Beth had nothing new to report.


Discussion of telescope/instruments report:

Mark said he thought things were going well, overall.  There was a problem 
with a Hub process memory leak, which was solved.  The weather has been 
exceptionally warm and clear of late.  DIS is running well, with tuning of 
the cryotigers and set points in progress.  Agile has been well-behaved, 
and SPIcam is back on the sky with its new vacuum window.  SPIcam's recent 
problem with noise pickup appears to have been solved with a new 
electrical isolation disk.

Mark Klaene's detailed report for the reporting period follows:


 	      3.5-m Telescope and Instruments Highlights
 	                11/01/12 through 11/27/12

0) Overview

Productive period with good seeing at times and some weather-associated 
lost time.  One occurrence of loss time due to a HUB memory leak from a 
test program that was running.

1) Telescope

The telescope performance has been nominal this period.

2) Instruments --

DIS has been operating nominally.  Both red and blue cameras have been 
functioning normally with good vacuum and temperature.  Adjustments were 
made to the cryotiger systems that affected blue temperature slightly. 
We are seeing slight variations in CCD temps with diurnal variations of 
ambient temperature as well as seasonal variations.  The diurnal 
variations vary with cameras to 2-6 deg K but often do not settle out 
until 2 to 4 hours after sunset.  We are looking at making improvements 
and talking with the cryotiger manufacturer about options.

Agile had no issues this period.

SPICam window was replaced.  The AR coatings on the new window are likely 
better than the old window (we had no data on the old window).  The new 
window coating plots will be added to the documentation.  A new aluminum 
nitride disc was manufactured and installed to provide electrical 
isolation between the cold strap and the cyrotiger coldhead.  Bias levels 
are comparable to pre-shutdown levels.  SPIcam is now back in routine 
operation (as of Nov 19).

The new TripleSpec cart is in routine operation.  We had one failure of 
the cart's jacking mechanism that has been repaired.

Echelle is operational.  We are testing new cables this week that should 
improve flip mirror operations.  The slitviewer camera had another shutter 
failure.  This is a design flaw in the APOGEE camera.  We are now running 
on Ed Leon's in-house driver circuit indefinitely.

NIC-FPS is operating in shared-risk observing mode due to occasional image 
corruption issues.


GIFS status and training opportunity:

Suzanne discussed the status and plans for GIFS, the "Goddard 
Integral-Field Spectrograph", formerly known as GFP.  Mike McElwain (the 
new PI at Goddard following Bruce Woodgate's retirement - thanks Bruce for 
all your efforts over the years!) and Carol Grady wish to train users in 
its use.  Users were asked to send observing proposals, although not much 
feedback has been so far received.  Bill mentioned that GIFS is usable as 
a direct imaging instrument, without the coronal stop, in the wavelength 
range from 4800A to 7800A, which hopefully will soon be increased to 4000A 
to 9000A.  The integral-field spectrograph has about two-thirds of its 
functionality implemented -- the high-resolution grating is not yet 
available, and a draft version of the GIFS users manual has just been 
posted on the APO website.  The Fabry-Perot option is not currently 
available for users.  Carol Grady has also offered to help users learn how 
to run the data reduction software.  Suzanne said that the on-site 
training will likely be in Q2, to insure that the instrument software, 
manuals, etc. are fully ready.  Several of the users committee agreed that 
it was better to wait a couple months than to try to do the training 
before everything is in place and properly working.  Bill said that an 
announcement will be sent to apo35-general before the next proposal cycle. 
Suzanne said that the training will be done during DD time, and then the 
science programs will be block-scheduled adjacent in time to the training. 
Users are encouraged to review the users manual, and to contact Mike and 
Carol for more information if they wish to put in a proposal for Q2 and to 
attend the on-site training.


DIS slitviewer:

The slitviewer focus mechanism was upgraded in November and appears to be 
working well.  Mark stated that the DIS slitviewer upgrade is making good 
progress.  We are planning to replace the camera and working on an optics 
upgrade.  Bill added that the current system has a 2-1/2 arcminute FOV on 
a side, and we are aiming to get up to 4 arcminutes with the new system. 
Also, the slit-view data should be less noisy.


Issue with image headers (time reporting):

Bill reported that we have been inconsistent in the header UT1 and LST 
time stamps since the summer shutdown.  This was fixed in early November, 
but users need to know that the header times were incorrect, often by a 
large amount.  In any case, Suzanne recommended that users should try to 
use TAI.


Echelle update:

Jon Holtzman said he is still working on the echelle throughput question, 
i.e., did it improve as hoped for after the new grating installation last 
summer.  The instrument is fully operational, however.  Mark mentioned 
that the echelle slitview shutter failed again -- the problem is with the 
Apogee electronics, and we are now using a home-brew driver circuit that 
seems to be working well.


TUI support ending for Mac OS 10.5:

Suzanne reminded us that TUI support for Mac OS 10.5 will be ending in 
June 2013 when summer shutdown starts.  All users are encouraged to 
upgrade to at least Mac OS 10.6 (the current version if Mac OS 10.8).


Other business:

Mark said that although we continue observing operations through the 
holidays, between Dec. 22 and Jan. 1 the day staff support will be 
minimal.  E-mail and phones will continue to be monitored as usual.



[open from previous meeting]:

===> ACTION:  Send new ARC partner suggestions to Suzanne and Rene (before 
next meeting).

 	Status:  Open, committee members need to keep sending suggestions. 
We need the names of institutions (of any size) and a named contact. 
Send info to Rene, Suzanne and Bruce.

===> ACTION:  Users committee members should discuss APOGEE fiber feed to 
3.5-m proposal with their users and forward feedback, and ideally name(s) 
of interested users who would be willing to participate in an ATI proposal 
to Suzanne, Mark and Fred Hearty.

 	Status:  Open, see notes from previous meeting.  Suzanne didn't 
hear much feedback.  At the ARC Board meeting, the possibility of keeping 
the APOGEE spectrograph at APO hinges on the success of getting funding, 
possibly through the MRI grants program, to build a 2nd spectrograph for 


Next meeting:  The next Users Committee phonecon will be on Monday 15 
January at 8:30 AM Pacific Time.  The agenda and other materials will be 
sent to the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1298 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to